OpenStudy (anonymous):

Two urns contain white balls and yellow balls. The first urn contains 3 white balls and 6 yellow balls and the second urn contains 3 white balls and 8 yellow balls. A ball is drawn at random from each urn. What is the probability that both balls are white?

OpenStudy (anonymous):

wats up

OpenStudy (anonymous):

if \(W_1\text{ and } W_2\) are events of getting a white ball from urns 1 and 2, respectively, then \(P(W_1 \text{ and } W_2)=P(W_1) \cdot P(W_2) \)

OpenStudy (anonymous):

3/10?

OpenStudy (anonymous):

\[P(W_1)=\frac{3}{9},\;P(W_2)=\frac{3}{11}\]\[P(W_1\text{ and } W_2)=\frac{3}{9}\cdot \frac{3}{11}=\frac{1}{11}\]

OpenStudy (anonymous):

no 2/23?

OpenStudy (anonymous):

33**

OpenStudy (anonymous):

no dear, it's not addition. it's multiplication.

OpenStudy (anonymous):

well those r what i have to pick from .... 3/33, 2/10,1/9,1/11

OpenStudy (anonymous):

the last choice is 1/11

OpenStudy (anonymous):

thank u